OpenStudy (anonymous):

how to simplify 6 and 1/4 times 4 and 2/3?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\[6\frac{ 1 }{4 }\times4\frac{ 2 }{ 3}\]

OpenStudy (anonymous):

you mean that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(6•4+1)/4=25/4 (4•3+2)/3=14/3 (14•25)/(4•3)=\[29\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (tkhunny):

I like this... (6 + 1/4)*(4 + 2/3) = 24 + 4 + 1 + 1/6 = 29 1/6 Just like any other "binomials".